# Jonathan Light1  Min Cai2  Weiqin Chen1  Guanzhi Wang5  Xiusi Chen3 Wei Cheng4  Yisong Yue5  Ziniu Hu5 1

### v1

@@ 3 Experiments
Figure 3: Left: Example performance of value heuristics strategy tree for GOPS. Points represent the final evaluation score and generation of improved functions. Lines depict which function the improved function evolved from. Our method is able to escape local maxima and continue exploring the strategy space. Note that these are final game-play scores against a fixed distribution of opponents, which differs from the intermediate self-play scores used to select the best strategy to improve upon during the improvement process. Right: Comparison of different improvement methods on 6 card GOPS, where we play 9 functions generated by each method against each other for 1024 games total.
We demonstrate the effectiveness of our self-improvement process through experiments against different improvement benchmarks. We tested our method on (1) GOPS, a two-player zero-sum card game (see B for rules) and (2) Avalon, a five or more player team-based discussion game (see A for rules). For Avalon dialogue generation, we specifically benchmark on the Merlin role and Assassin roles, representing both Good and Evil sides. Experiments were run with gpt-3.5-turbo-0125, with more ablation studies in Appendix J. Details on how we implemented policies from learned strategies and acquired feedback are described in Section 2.1, 2.3, and 2.4, and Appendix D, E, and F.
@@ 3.1 Different LLM Improvement Methods
Our results are shown in Table 1, where the method of collecting feedback (simulational self-play) is the same but we vary the improvement method. The number of new strategies generated by each method is also held constant. The gameplay scores of playing the strategies generated by each method against each other on GOPS is shown in Figure 3 right. Even when we control for the number of output tokens generated by the LLM, we see that our method still achieves higher performance as shown in Figure 4. We believe that the reason why we see higher performance in our method is because (1) the idea queue helps test which incremental improvements are helpful and guide the search process and (2) our strategy and idea selection policy help us explore the strategy space more efficiently and escape local maxima as shown in Figure 3 left.
Table 1: Comparison of different self-improvement methods. The improvement process was run with the same number of strategies generated for each method (40 for GOPS, 24 for Avalon) on the same seed functions, all with GPT3.5. We collect 9 functions from each process and play them against each other (a total of $5\times 9=40$ different agents), reporting the average number of points you win over your opponents for GOPS and the average winrate for Avalon. The median scores ($\pm$ IQR/2) for each method are shown. For the dialogue guide, we show the improvement over the baseline seed function, which has a score $z=-0.875\in[-2,2]$, a rating from opponents which we describe in F.
@@
3.2 LLM-improvement vs. Reinforcement Learning (RL) Training
@@
GOPS Value Heuristic $0.72\pm 0.00$ $0.31\pm 0.35$ $0.66\pm 0.60$
Avalon Value Heuristic $0.23\pm 0.00$ $0.30\pm 0.11$ $0.38\pm 0.12$
3.3 Feedback Quality and Reward Signal
Recent works tend to focus on either using another LLM to critique the generations and thus provide feedback  (Madaan et al., 2024; Shinn et al., 2024), observations from real environment interactions  (Nottingham et al., 2024), or some combination of the two  (Wang et al., 2023). Since our method learns how to internally simulate the opponents by learning a strategy and thus gain insights into the opponent’s policy, we are able to acquire better quality feedback than either approach. We benchmark our feedback acquisition method against (1) using a LLM-critic and (2) trajectory feedback from interactions against a fixed opponent policy. Our results are shown in table 3, where our method achieves better performance on both action planning in GOPS and dialogue generation.
@@ 3.3 Feedback Quality and Reward Signal
[…]
GOPS Value Heuristic $-0.27\pm 1.1$ $0.089\pm 0.86$ 0.87 ±$\pm 1.5$
Avalon Dialogue Guide $0.37\pm 0.063$ $0.62\pm 0.13$ 0.88 ±$\pm 0.063$
[…]
Method Setting Detailed Stats
total winrate (%) assass. acc (%) deduc. acc (%)
ReAct (Yao et al., 2023) w/o discussion 10 20 51
w/ discussion 10 20 70
[…]
w/ discussion 50 100 48
@@ 4 Related Work
AI in strategy games. AI has been applied to great success in board games. AlphaGo and MuZero demonstrated the power of combining MCTS, deep learning, and feedback generation using self-play in games such as Go, Chess, and Shogi Silver et al. (2017); Schrittwieser et al. (2020). Language models can also be trained on human in-game discussion data and integrated with another separately trained action planner to play board games with dialogue  † et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al. (FAIR). We build upon the AI for games literature by showing that LLMs can accomplish both (1) the training of a value heuristic like that in AlphaGo through self-play more efficiently than RL and (2) dialogue generation in discussion games with no human examples. These adversarial environments are not just limited to board games. For example, there has been recent interest on creating LLM-agents that can negotiate Abdelnabi et al. (2023); Fu et al. (2023), which our method can also be applied to.
5 Limitations and Conclusion
@@ Appendix B Game of Pure Strategy (GOPS) Game Description
Appendix C Improvement Process Implementation Details
@@
Appendix D Avalon Agent Implementation Details
@@
D.1 Dialogue Analyzer (Discriminator)
@@
We show examples of such prompts in Appendix I.
D.2 Action Planner
@@
More specifically, in our search implementation, at the first layer, we first sample across possible hidden states $\boldsymbol{s}\sim\widehat{\boldsymbol{b}}$ according to the belief prior. At the second layer (i.e. the first action stage $t$), we calculate expected $q$-values for each action $a\in\mathcal{A}$ that the main player can take if the other players play actions $\boldsymbol{a}\sim\widehat{\boldsymbol{\Pi}}_{t}$ according to the predicted joint distribution. In subsequent action stages, the search process will assume that other players play according to their policy simulated and induced by the value heuristic that is not dialogue dependent. We then take the best response action $a^{*}_{i}=\mathop{max}(\boldsymbol{\pi}^{i})$ as the intended action. Since this is a partial information game, expected $q$-values are taken across information sets, not states. We describe how our action planner is implemented in more detail in Appendix E.
D.3 Dialogue Generation
@@
We show examples of such prompts in Appendix I.
Appendix E Value Heuristic Implementation Details
@@
Appendix F Dialogue Guide Improvement Evaluation Implementation Details
@@
We first generate a synthetic dataset by simulating a game of Avalon with initial dialogue and move policies $\boldsymbol{\phi}$. Given the dialogue guide $\sigma$ we want to evaluate, we then sample ‘scenarios’ from the dataset. A scenario consists of a game state, intended action, and private information in the simulated trajectory. We create an Avalon agent like the one we described in D for each player in the game, initialized with their corresponding private information. The Avalon agent is then asked to generate dialogue using the dialogue guide $\sigma$.
@@
The dialogue analyzer (discriminator) is described in more detail in Appendix D and the specific generation and analysis prompts are shown in Appendix I.
Appendix G Value Heuristic LLM Prompt and Output Examples
G.1 System Prompts
@@
G.2 Idea Generation Examples
@@
G.3 Strategy Implementation Examples
@@ G.3.4 Example 4
Appendix H Dialogue Guide LLM Prompt and Output Examples
@@ H.1 System Prompts
H.2 Idea Generation Examples
@@
H.3 Strategy Implementation Examples
@@ H.3.3 Example 3
Appendix I Dialogue Generation and Analysis Example LLM Inputs and Outputs
@@ I.3 Dialogue Analysis
Appendix J Other Ablation Studies
J.0.1 Base LLM Model
@@
Appendix K Details on Learning the Value Heuristic via Reinforcement Learning
@@
Appendix L Experimental Compute Resources
@@
Appendix M Improvement method baselines
Line search always reflects and improves upon the latest improved strategy. In other words, we implement a strategy, collect feedback from the environment on the strategy, then improve it using the LLM, rinse and repeat, in Reflexion style  (Madaan et al., 2024).
Greedy search selects the best strategy from the last generation of improved strategies to improve upon each improvement cycle. Each improvement cycle, we take the best strategy from the previous improvement cycle and create $n$ different improved versions of it. We then collect feedback from the environment on these $n$ new strategies, and pick the best one to progress to the next improvement cycle. This is how Eureka  (Ma et al., 2023) improves reward functions using LLMs, similar to an evolutionary algorithm
Best first search improves upon the $k$ best strategies generated in any iteration of each improvement cycle, similar to beam search. However, unlike beam search, we can set the branching factor to greater than one, where we generate more than one strategy from each previous strategy selected. Tree of thought explored a similar approach for self improvement  (Yao et al., 2024).
Best first search with thought asks the LLM to improve upon the thoughts used to generate the $k$ best strategies before improving the strategy itself while conducting best first search. This is similar to adding a React style thought and action to the improvement step  (Yao et al., 2022b).
@@
is our method that uses an additional idea queue $Q$ and an idea generation step to guide the improvement process.
All methods were run with the same budget, i.e. generating the same number of improved strategies, in our experiments where we compared the methods.

### v2

@@ 3 Experiments
Figure 3: Left: Example performance of value heuristics strategy tree for GOPS. Points indicate the final evaluation scores and the generation of improved functions, with lines showing the evolutionary path between functions. Our method successfully escapes local maxima and continues exploring the strategy space. These scores reflect final gameplay results against a fixed set of opponents, differing from the intermediate self-play scores used to guide strategy improvements. Right: Comparison of different improvement methods on 6-card GOPS, where 9 functions generated by each method play against one another over 1024 total games.
[…]
@@ 3.1 Different LLM Improvement Methods
Our results, presented in Table 1, use the same feedback method (simulation-based population self-play) while varying the improvement methods, with a constant number of new strategies generated across all methods. Figure 3 (right) shows the gameplay scores of these strategies on GOPS. Even when controlling for the number of output tokens generated by the LLM, our method outperforms the others, as demonstrated in Figure 15. We attribute this to (1) the idea queue, which helps test incremental improvements and guide the search process, and (2) our strategy and idea selection policy, which more efficiently explores the strategy space and escapes local maxima (Figure 3 left).
Table 1: Comparison of different self-improvement methods. The improvement process was run with the same number of strategies generated for each method (40 for GOPS, 24 for Avalon) on the same seed functions, all with GPT3.5. We collect 9 functions from each process and play them against each other (a total of $5\times 9=40$ different agents), reporting the average number of points you win over your opponents for GOPS and the average winrate for Avalon. The median scores ($\pm$ IQR/2) for each method are shown. For the dialogue guide, we show the improvement over the baseline seed function, which has a score $z=-0.875\in[-2,2]$, a rating from opponents which we describe in G.
@@
3.2 Scaling Inference-Compute can Continually Improve Strategy
[…]
3.3 LLM-improvement vs. Reinforcement Learning (RL) based methods
We demonstrate the effectiveness of our method compared to traditional RL-based approaches for learning a good policy. Specifically, we show that our method learns a value heuristic function more efficiently than deep RL, as used in AlphaGo, MuZero  (Silver et al., 2017; Schrittwieser et al., 2020), and DeepRole  (Serrino et al., 2019). While AlphaGo uses MCTS  (Kocsis & Szepesvári, 2006) with a deep value network, DeepRole combines counterfactual regret minimization with a deep value network  (Moravčík et al., 2017; Zinkevich et al., 2007). We adapt both algorithms to our setting.
[…]
Table 2: Comparison of reinforcement learning vs our method. We run each process 10 times, taking the best strategy generated by each run and playing them against each other for 1024 games. Average win-rates, total number of self-play episodes, and training transition steps across best generated strategies are shown here.
Setting Metric VS Alpha-go VS DeepRole
[…]
GOPS Winrate $-0.39\pm 0.71$ $0.33\pm 1.21$ $-0.56\pm 0.57$ $1.14\pm 0.30$
Self-play eps. 320 episodes 320 episodes 320 episodes 320 episodes
Trans. steps $\sim 3,840$ steps 100 steps $\sim 3,840$ steps 100 steps
Avalon Winrate $0.30\pm 0.11$ $0.38\pm 0.11$ $0.33\pm 0.12$ $0.44\pm 0.08$
Self-play eps. 160 episodes 160 episodes 160 episodes 160 episodes
Trans. steps $\sim 24,000$ steps $60$ steps $\sim 24,000$ steps $60$ steps
3.4 Feedback Quality and Reward Signal
We benchmark our feedback method (population based self-play) against (1) an LLM-critic  (Madaan et al., 2024) and (2) trajectory feedback from a fixed opponent policy, with results in Table 3 showing superior performance in both GOPS action planning and dialogue generation. This shows the effectiveness of our evolutionary population based self-play approach, and that of our strategy abstraction that learns a strategic profile for all players.
[…]
Setting/Method Metric LLM-critic Fixed opponent Population-based Self-play
GOPS Performance $-0.27\pm 1.1$ $0.089\pm 0.86$ 0.87 ±$\pm 1.5$
# of opponents 0 1 4
Avalon Performance $0.37\pm 0.063$ $0.62\pm 0.13$ 0.88 ±$\pm 0.06$
# of opponents 0 1 4
@@ 3.4 Feedback Quality and Reward Signal
Metric VS ReAct VS ReCon
@@
Winrate 47.5 $\pm$ 2.5 52.5 $\pm$ 2.5 38.9 $\pm$ 5.5 61.1 $\pm$ 5.5
#Tokens per round 56 $\pm$ 14.3 164.3 $\pm$ 27.7 245.7 $\pm$ 21.2 248.2 $\pm$ 24.1
@@ 4 Related Work
AI in strategy games. AI has been applied to great success in board games. AlphaGo and MuZero demonstrated the power of combining MCTS, deep learning, and feedback generation using self-play in games such as Go, Chess, and Shogi (Silver et al., 2017; Schrittwieser et al., 2020). Language models can also be trained on human in-game discussion data and integrated with another separately trained action planner to play board games with dialogue  († et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR). We build upon the AI for games literature by showing that LLMs can accomplish both (1) the training of a value heuristic like that in AlphaGo through self-play more efficiently than RL and (2) dialogue generation in discussion games with no human examples. These adversarial environments are not just limited to board games. For example, there has been recent interest on creating LLM-agents that can negotiate (Abdelnabi et al., 2023; Fu et al., 2023), which our method can also be applied to.
5 Conclusion
@@
[…]
Furthermore, our results suggest broader implications for the development of LLM-driven decision-making agents across a range of complex domains. By embedding flexible and adaptive learning mechanisms, our framework paves the way for more autonomous systems capable of mastering tasks with minimal human intervention. This opens new avenues for the application of LLMs in areas such as multi-agent systems, reinforcement learning, and general AI, where strategic planning and self-improvement are key.
@@ Appendix B Game of Pure Strategy (GOPS) Game Description
Appendix C Limitations
Appendix D Improvement Process Implementation Details
Figure 9: Overview of our high-level strategy improvement method: The process alternates between two steps—idea generation and strategy improvement.
@@ Appendix D Improvement Process Implementation Details
Appendix E Avalon Agent Implementation Details
@@
E.1 Dialogue Analyzer (Discriminator)
@@
We show examples of such prompts in Appendix J.
E.2 Action Planner
@@
More specifically, in our search implementation, at the first layer, we first sample across possible hidden states $\boldsymbol{s}\sim\widehat{\boldsymbol{b}}$ according to the belief prior. At the second layer (i.e. the first action stage $t$), we calculate expected $q$-values for each action $a\in\mathcal{A}$ that the main player can take if the other players play actions $\boldsymbol{a}\sim\widehat{\boldsymbol{\Pi}}_{t}$ according to the predicted joint distribution. In subsequent action stages, the search process will assume that other players play according to their policy simulated and induced by the value heuristic that is not dialogue dependent. We then take the best response action $a^{*}_{i}=\mathop{max}(\boldsymbol{\pi}^{i})$ as the intended action. Since this is a partial information game, expected $q$-values are taken across information sets, not states. We describe how our action planner is implemented in more detail in Appendix F.
E.3 Dialogue Generation
@@
We show examples of such prompts in Appendix J.
Appendix F Value Heuristic Implementation Details
@@
Appendix G Dialogue Guide Improvement Evaluation Implementation Details
@@
We first generate a synthetic dataset by simulating a game of Avalon with initial dialogue and move policies $\boldsymbol{\phi}$. Given the dialogue guide $\sigma$ we want to evaluate, we then sample ‘scenarios’ from the dataset. A scenario consists of a game state, intended action, and private information in the simulated trajectory. We create an Avalon agent like the one we described in E for each player in the game, initialized with their corresponding private information. The Avalon agent is then asked to generate dialogue using the dialogue guide $\sigma$.
@@
The dialogue analyzer (discriminator) is described in more detail in Appendix E and the specific generation and analysis prompts are shown in Appendix J.
Appendix H Value Heuristic LLM Prompt and Output Examples
H.1 System Prompts
@@
H.2 Idea Generation Examples
@@
H.3 Strategy Implementation Examples
@@ H.3.4 Example 4
Appendix I Dialogue Guide LLM Prompt and Output Examples
@@ I.1 System Prompts
I.2 Idea Generation Examples
@@
I.3 Strategy Implementation Examples
@@ I.3.3 Example 3
Appendix J Dialogue Generation and Analysis Example LLM Inputs and Outputs
@@
Appendix K Other Ablation Studies
K.0.1 Base LLM Model
@@
Appendix L Details on Learning the Value Heuristic via Reinforcement Learning
@@
Appendix M Experimental Compute Resources
@@
Appendix N Improvement method baselines
Line search is a straightforward iterative process that continuously builds upon the most recent improved strategy. In each iteration, a strategy is generated and evaluated based on feedback from the environment. The feedback is then used to enhance the strategy via the LLM, and this cycle is repeated—similar to the Reflexion framework (Madaan et al., 2024). The essence of line search lies in its focus on immediate feedback, where only the latest strategy is considered for improvement, ensuring that progress is always aligned with the most recent understanding of the problem space.
Greedy search takes a more competitive approach by selecting the best-performing strategy from the last generation to serve as the foundation for the next cycle of improvements. During each cycle, the best strategy is used to generate $n$ variations, each an attempt at improvement. These variations are evaluated, and feedback from the environment is gathered to determine the best candidate for the subsequent cycle. This method is inspired by the Eureka framework (Ma et al., 2023), where LLMs are leveraged to iteratively refine reward functions in a manner akin to evolutionary algorithms, focusing on a winner-takes-all selection mechanism.
Best-first search generalizes the improvement process by considering the top $k$ strategies at each iteration, as opposed to focusing on a single best option. This is reminiscent of beam search but with a flexible branching factor that allows for the generation of multiple strategies from each selected candidate. By expanding multiple promising pathways simultaneously, the search can escape local optima and explore a broader solution space. This approach is related to the Tree of Thought method (Yao et al., 2024), which similarly employs a branching mechanism to explore various strategies in parallel during self-improvement cycles.
Best-first search with thought enhances best-first search by incorporating a reasoning layer into the improvement process. Before generating new strategies, the LLM is prompted to first refine the underlying “thoughts” or decision-making processes that led to the generation of the top $k$ strategies, including possible improvement ideas, before generating a new strategy. This meta-cognitive step draws from the React framework (Yao et al., 2022b), which emphasizes reasoning and actions during strategy development, adding an introspective element that encourages deeper exploration of strategy refinements.
@@
[…]
All methods were run with the same computational budget—i.e., each method was allowed to generate the same number of improved strategies—ensuring a fair comparison across approaches in our experiments.
We display a scaling curve for the various improvement methods in Figure 15.